Deployment team

→ Europe/Zurich
EVO - GridPP Deployment team meeting

EVO - GridPP Deployment team meeting

Jeremy Coles
Description
- This is the weekly DTEAM meeting - The intention is to run the meeting in EVO: http://evo.caltech.edu/evoGate/. Join the meeting in the GridPP Community area - The phone bridge number is +41 22 76 71400. Th phone bridge ID is 344539 with code: 4880
Present Andrew Elwell Graeme Stewart Pete Gronbech Jeremy Coles Brian Davies Raja Nandakumar Frederic Brochu Duncan Rand Steven Burke Mingchao Ma Derek Ross Greig Cowan Jens Jensen Dave Colling [part] Experiment status ---------------------- LHCb - Raja CCRC08: going well files into ral, reconstruction OK using castor old reconstruction in parallell - staging in off dcache - seems slow - big backlog - talking to Derek and Tim to resolve. CMS - Dave C CCRC - going reasonably well - most sites doing well - some sites have minor issues which are being dealt with. seeing 0.1 to 0.2% failure on jobs so far. (from mail - 28010 jobs over 21 siteas with a total of 71 failures) larger scale Multiuser analyis due to start early june - only a few people doing analysis just now. Greig - do you have an update on imperial dCache? - No. Duncan doesn't know either. Acknowledge its flakey - seeing timeouts on gridFTP managed by FTS. bugs open with developers. Mona and Kostas following up. ATLAS - Graeme production was going well last week until Sat when UK DDM box locked up. (known bugs) - restored on Sat PM then RAL Oracle failed and FTS locked up. discovered ~10pm at night. then discovered credential delegattion had got corrupted. Switched from Marios to Kors' certificates for transport within UK - went sorta OK. Major outage fixed out of hours - process now works - good improvement. one falloout thats affecting *some* T2s with FTS - delegated credentials missing on the channels. Brunel, ECDF, lancaster, ucl-hep. manchester, oxford. -- sems to be SRMv1 endpoints affected. Spoken to Matt to switch any transfer involving a space token to use SRM2 - may be that if it uses SRM1 then they havent been passed source space token? ongoing.... (no space token used on GET) may be issue wiuth URL copy channel from 1->2? Production therefore only at ~50% of UK BUT they got a whole pile of bad tasks (4000 jobs) into Q - expect 3* that failures - should be killed officially by atlas MC coordinator rather than Graeme (just to let sites know) CCRC - T1-T1 transfers started today - RAL looks good. 98% as dest ~20 errors of src timeouts (SARA / FZK / Taiwan) RAL as source - Mostly ok apart from CNAF and FZK. [aside from Derek - They had already started working on the nagios alert before GS's RT ticket hit] Greig - Gridview not showing any atlas data but it's there in the dashboard. - perhaps some lag?] ... discussion ensued about weaknesses of gridview - esp wrt to dcache parsing April reliability figures: Scotgrid poor as we were closing old edinburgh. (also had poor reliability) ECDF getting better. (last week bad again) LT2 - QMUL and UCL-Cent pulled figures down. Brunel poor too - Still missing a member of staff. Henry doing in spare time (as is Duncan as coord) Northgrid - Liverpool in feb bad, otherwise back up to 93% Southgrid - Oxford upgraded in April + migrating nodes between sites. RAL dcache issues and JET - unknown. -- but Accounting in April = matched scotgrid :-) Storage Avail plots - http://www.gridpp.ac.uk/wiki/GridPP_storage_availability_monitoring ANALYSIS WORKSHOP --------------------------- -- not going ahead but a larger user session at gridpp21 mtg -- expansion of UIs (say onto laptops etc) * Rather than full UI on local machine (need an SL4 VM ideally) * offer UI at sites that users can GSISSH into <<-- recommendatiion * tarball install available on all group machines in depts. <<-- recommendation * need CRL / Configs - not trivial task Most people submit from CERN - more now starting in UK from Site breakdown - 3 at lancs, how many are needed? what should we offer - one per site, one per T2 - realistically one (or more) per site depending on install method for the UI you need a *real* account not a pool a/c - may be issue with site admins? Team update: - ------------------ Pete G - HEPIX last week. benchmark talks good - quad core AMD chips _can_ perform better than intel but v close. (don't panic if you got intel) - SL update - xfs will go into vanilla SL5 and in contribs on SL4. Security patches for SL309 for consiiderable time (oct 2009/2010?). Oxford work - Older nodes on SL4. New torque server for all 3 CEs. Greig helping with DPM. Need to move 1/2 cluster up to new site. Andrew - Nagios / WLCG stuff. DPM usage and deliverables for next funding Duncan @brunel CCRC08, getting CMS sam tests passed. bandwith issues etc. @imperial - Mona installed tarball - getting atlas jobs running @rhul - maui issues - working with vendor @QM - atlas install issues last week - coulnt install as atlassgm. @ucl-cent - VOMS issues. CE waiting for network ports to be opened Steven B - Docs / User support. userguide supposed gto be taken over. Gridpp website maintenance Been working on glue schema (glue2.0) - public comment from monday Mingchao - security workshop in oxford last week Derek R - new CE for atlas (2 more this week one CMS one LHCb) - should GS use ce03? - waiting to watch load. watching CCRC08 dCache issues - stageoff issues - ongoing Greig - HEPIX working group - benchmarking f/s and storage tech. Working on LHCb mc data storage at edinburgh + grif. (slight change to computing model). SLC vs SL4? had probs getting jobs running on ECDF. official line was we don't support non-slc distros. [RAJA - who said this? email exchange -- they consider SL / SLC equal] - seems like they were wrong. Manpower forcussed on dirac3 development not dirac2 support. no prodn jobs in Q anyway. hopefully will start prodn soon on dirac3 Graeme - atlas + procurement. Formal kickoff starts tomorrow. Debugging ECDF problems with gpfs. (load on gpfs daemons sucking CPU) - 2 of the 4 sites that had atlas s/w on gpfs space have abandoned it. seems to happen at all points in the job - doesn't make much sense.- will be running some standalone jobs to test on gpfs / local disk. Durham - all OK and quiet. Jens - updating venezuala CA. new castor info provider now live (docs to follow). busy with docs for allhands, OGF, pma etc etc. TICKET UPDATES --------------------- 28457 - Jeremy to follow up 33373 - Raja thinks exchange of emails - roberto thinks it can be closed. 32029 - old site. Greig to investigate. ... more next week AOB ----- GDB tomorrow - Alessandra volunteered to go 1st representing T2s Agenda at http://indico.cern.ch/conferenceDisplay.py?confId=20229 Peter Love to attend Dteam as lack of info about say CCRC progress? -- can attend the daily CCRC specific calls. -- JC will discuss with alessandra for T2 dissemination. -- reluctance to turn dteam into tb-support meeting All Hands - anyone submitting? (~4 from gridpp) focus has changed - more about what users can do? some of the accepted ones are more user-specifiic. hepsysman - meeting mid June at RAL? (post CCRC and last minute before data starts) GS and AE were volunteered. Thanks...
There are minutes attached to this event. Show them.
    • 11:00 → 11:20
      Experiment problems/issues 20m
      Review of weekly issues by experiment/VO - LHCb - CMS - ATLAS - Other
    • 11:20 → 11:35
      Site availability review 15m
      - Recent SAM performance - http://hepwww.ph.qmul.ac.uk/~lloyd/gridpp/samtest.html - Storage availability
      April WLCG T2 reliability
    • 11:35 → 11:45
      Improving the availability/number of UIs 10m
      - Is the documentation for deployment up to date? - Are there still reservations about users installing locally - How can we make more UIs available
    • 11:45 → 11:55
      Round the team update 10m
      - Brief summary of current activities
    • 11:55 → 12:05
      Tickets/action review 10m
      - https://gus.fzk.de/download/escalationreports/roc/html/20080513_EscalationReport_ROCs.html
    • 12:05 → 12:10
      AOB 5m
      - There is a GDB tomorrow: http://indico.cern.ch/conferenceDisplay.py?confId=20229.